# Linguistics and Human Brain: A Perspective of Computational Neuroscience

### v1

@@ Abstract
Elucidating the language-brain relationship requires bridging the methodological gap between linguistics’ abstract theoretical frameworks and neuroscience’s empirical neural data. As an interdisciplinary cornerstone, computational neuroscience formalizes language’s hierarchical and dynamic structures into testable neural representation models via modeling, simulation, and data analysis, enabling computational dialogue between linguistic hypotheses and neural mechanisms. Recent advances in deep learning, particularly large language models (LLMs), have further advanced this inquiry: their high-dimensional representational spaces offer a novel scale to explore the neuroal basis of linguistic processing, while the “model-brain alignment” framework provides a methodology for evaluating the biological plausibility of language-related theories.
This review synthesizes interdisciplinary progress from a computational neuroscience perspective. First, it outlines the core connotations of classical linguistic frameworks (generative grammar, functional linguistics, cognitive linguistics), their cross-cultural evolutionary characteristics, and key challenges hindering neural alignment—including insufficient quantitative mechanisms, abstract construct inaccessibility to neural measures, and limited treatment of dynamics/plasticity. Second, it details the methodological foundation of linguistics-neuroscience dialogue, focusing on four technical pillars: neural activity measurement (fMRI, EEG, MEG, fNIRS, ECoG, SEEG, etc.), linguistic numerical representation, language model evolution (from statistical models to LLMs), and neural coding frameworks linking model representations to brain signals, illustrated by a model-brain alignment case study. Third, it summarizes key findings: early computational models’ insights into predictability and structural properties, and LLM-driven neural alignment progress across cross-modal interaction, inter-brain coupling, hierarchical computation, learning strategy sensitivity, and language plasticity. Finally, the review analyzes limitations of current research—including functional alignment without structural homology, technical constraints on real-time verification, biased research coverage, and narrow evaluation frameworks—and proposes future directions: enhancing model biological plausibility via spiking neural network-based brain-inspired LLMs, constructing cognitive-level alignment frameworks integrating memory/causality/metacognition, and expanding clinical translations (language disorder intervention, brain-computer interface upgrades).
@@ 1 Introduction
Recent advances in artificial intelligence, particularly in deep learning and Large Language Models (LLMs), have accelerated this integration further. Researchers can now explore linguistic structures and their correspondence with brain networks within higher-dimensional representational spaces [226, 214, 247]. LLMs provide computational platforms for examining classical linguistic phenomena such as semantic integration, long-distance dependencies, and predictive processing. They also support the emerging framework of model–brain alignment, wherein internal model representations are used to predict and explain neural responses during language processing [56, 5, 117]. This development marks a shift from early conceptual modeling toward quantitative mappings between large-scale linguistic data and high-resolution neural signals.
This paper provides a systematic review of how computational neuroscience serves as a methodological bridge between linguistics and neuroscience. Section 2 surveys classical linguistic frameworks and their development across languages, identifying the core challenges that have historically hindered their alignment with neural data. Section 3 introduces key methodological foundations, including neural data acquisition, word embeddings, modern language models, and neural coding frameworks. Section 4 reviews applications of classical computational neuroscience models to language comprehension, followed by a synthesis of recent progress in LLM-driven neural alignment, covering domains including cross-modal representation, cross-brain coupling, hierarchical computation, and learning mechanisms. Section 5 analyzes current methodological limitations and theoretical challenges, and outlines promising future research directions, such as enhancing biological plausibility, developing cognitive-level alignment frameworks, and expanding clinical and brain–computer interface applications. Finally, the paper summarizes the main conclusions and discusses the long-term significance of computational neuroscience for fostering a deeper integration of linguistics and neuroscience.
@@ 2 Theoretical Foundations of Linguistics
Language has long been the central subject of study linking linguistics and neuroscience. Over time, linguistic theories have continuously developed through theoretical modeling, cross-cultural adaptation, and empirical investigation. At the same time, they have gradually revealed structural limitations that hinder direct alignment with neural mechanisms of the brain. Classical linguistic theories provide systematic frameworks for describing linguistic structure and function. However, when confronted with cross-linguistic diversity, dynamic language use, and evidence from neural implementation, their explanatory power remains limited [70, 60, 162, 85]. Therefore, in this section, we review the major linguistic traditions, their evolution across cultural contexts, and the key challenges that currently motivate interdisciplinary integration, which provides the theoretical background for later discussions on the necessity of computational neuroscience as a bridging framework. Fig. 1 outlines the framework of this section.
@@
Figure 1: Classical Linguistic Theories and Challenges for Bridging Linguistics and Neuroscience.
2.1 Frameworks of Classical Linguistic Theories
@@
Modern linguistics has developed several influential theoretical frameworks. Among them, generative grammar, functional linguistics, and cognitive linguistics describe language structure and cognition from different perspectives.
@@
Generative grammar, most prominently associated with Chomsky, views linguistic competence as an innate component of the human cognitive system [35]. This framework emphasizes the existence of universal grammar, arguing that children can acquire complex grammatical structures despite limited linguistic input. This ability is taken as evidence for internally constrained grammatical principles. The distinction between deep structure and surface structure further illustrates how abstract syntactic relations are transformed into observable linguistic expressions [36, 109]. Later developments, such as the Minimalist Program, attempted to reduce linguistic theory to a small set of core operations, most notably the merge operation, in order to simplify theoretical assumptions and narrow the gap between linguistic theory and cognitive implementation [134]. Nevertheless, these highly abstract constructs remain difficult to map onto concrete patterns of neural activity, which has led to sustained debate in neurobiological research.
@@
In contrast, functional linguistics treats language primarily as a system for social communication. It emphasizes that linguistic form is shaped by communicative needs and contextual factors [237]. Systemic functional linguistics proposes that language fulfills multiple functions, including ideational, interpersonal, and textual functions. Linguistic choices are therefore closely linked to communicative goals and social relations [89]. Differences between spoken and written language, such as ellipsis in conversation and structural completeness in writing, are interpreted as outcomes of functional adaptation [65]. While functional linguistics has been highly influential in discourse analysis and cross-cultural communication studies, its strong emphasis on context and function makes strict formalization difficult. This limitation complicates efforts to connect functional explanations with neural mechanisms.
@@
Cognitive linguistics further rejects the idea that language constitutes an autonomous module. Instead, it argues that language is grounded in general cognitive processes, including categorization, metaphorical mapping, and mental simulation [130]. Conceptual metaphor theory suggests that abstract concepts are understood through concrete experience, and that grammatical structures reflect symbolic representations of conceptual relations [122]. Embodied cognition accounts further propose that language comprehension relies on sensorimotor experience and mental simulation [9]. Although this framework provides a unified view of language and cognition, many of its core constructs are difficult to operationalize and quantify. This limits direct comparison with neural data.
@@
Overall, these three theoretical traditions explain language from the perspectives of innate mechanisms, communicative function, and general cognition. However, they largely remain at an abstract theoretical level. A substantial gap persists between linguistic descriptions and the concrete implementation of language in the neural system. This gap poses a major challenge for subsequent computational and neural investigations.
@@ 2.2 Cross-Cultural Evolution of Linguistic Theories
The existence of thousands of languages worldwide reveals extensive diversity in linguistic structure and expression. This diversity has continuously driven revisions and extensions of linguistic theory, while also highlighting the difficulty of constructing a unified explanatory framework [171, 47].
@@
Functional linguistics approaches cross-linguistic variation by focusing on cultural practices and communicative conventions. Phenomena such as honorific systems, context-dependent expressions, and discourse-level differences are incorporated into the analytical framework [89]. For example, cultures differ in their emphasis on politeness, indirectness, and social hierarchy, which in turn shapes linguistic form [10]. Although this perspective helps explain cultural variation, its strong reliance on specific cultural contexts makes it difficult to extract universally applicable principles.
Cognitive linguistics seeks a balance between universal embodied experience and culturally specific conceptualization [130]. Many spatial and emotional metaphors are shared across languages, suggesting common cognitive constraints. At the same time, systems such as spatial reference frames, color categorization, and kinship terminology exhibit clear cultural specificity [121]. The coexistence of universality and diversity indicates that language is shaped by both biological constraints and cultural experience.
Cross-linguistic research thus shows that language structure reflects shared cognitive foundations as well as accumulated cultural history [39]. However, this dual origin makes it difficult to distinguish which linguistic properties arise from common neural mechanisms and which result from cultural learning. Evidence from cross-writing-system studies of developmental dyslexia further challenges the assumption of a universal neural basis for language-related processes, showing that reading impairment engages distinct brain regions in alphabetic versus logographic systems, and that the neural manifestation of dyslexia is shaped by cultural writing systems rather than being fully universal [220]. This ambiguity further complicates the integration of linguistic theory with neural models.
2.3 Challenges in Classical Linguistic Theories
With advances in neuroimaging techniques and computational methods, language research has increasingly moved beyond purely theoretical analysis toward empirical integration with neural data [87]. In this interdisciplinary context, traditional linguistic theories have revealed several structural challenges. These challenges limit direct links between linguistic constructs and neural mechanisms, while also motivating the development of new computational and neural modeling approaches [238]. Broadly speaking, these difficulties involve the lack of quantitative mechanisms, limited research scope, and poor integrability of findings.
@@ 2.3 Challenges in Classical Linguistic Theories
    Several interpretations have been proposed in the cognitive science literature. Some link the N400 to lexical predictability, suggesting that more predictable words elicit smaller responses. Others interpret it as reflecting semantic integration difficulty [157]. Yet these accounts largely remain conceptual. They do not specify how prediction and integration are implemented computationally within neural circuits, which limits precise theory–data correspondence.
@@
    Language research often relies on highly controlled experimental designs to ensure reliability and interpretability [240]. Such designs typically focus on isolated or local linguistic phenomena, using highly uniform stimuli. As a result, findings may not generalize to the complexity of natural language use.
    Meanwhile, there is a long-standing mismatch of research scales between linguistics and neuroscience [49]. Linguistic studies often examine fine-grained structural distinctions, whereas neuroscience focuses on broader functional organization across brain regions. Even in interdisciplinary work, experiments frequently rely on simplified or prototypical materials, which limits ecological validity and generalizability.
@@
    While controlled experiments help identify local effects, they also contribute to fragmented results. Individual studies usually target specific phenomena, yet language itself resists decomposition into a small set of independent units. Differences in task design, stimulus materials, and measurement techniques further hinder direct integration across studies. For instance, different types of linguistic stimuli may elicit distinct activation patterns or electrophysiological responses. However, aggregating such findings rarely yields a coherent account of how the brain incrementally constructs the meaning of a complete sentence [99].
@@
    Many central linguistic concepts, such as recursion, thematic roles, and conceptual metaphor, are highly abstract and lack clear operational definitions. This makes them difficult to translate into measurable neural variables [193]. Moreover, neural activity is distributed across networks rather than localized to single regions. This further complicates attempts to establish simple mappings between linguistic functions and specific neural indicators [62].
@@
    There is no consensus on whether cross-linguistic structural differences arise primarily from innate mechanisms or from cultural learning and experience. Divergent theoretical positions make it difficult for neural studies to determine whether observed processing differences reflect biological constraints or experiential adaptation. This uncertainty hinders the construction of a unified neurocognitive model of language.
@@
    Traditional linguistic theories are largely based on static sentence analysis and struggle to account for dynamic aspects of real-world communication. These include real-time prediction, contextual updating, and social interaction processes in naturalistic conditions [189, 188]. In addition, human language abilities show substantial plasticity during child acquisition, second language learning, and recovery after brain injury. Traditional frameworks lack unified explanations for these developmental and adaptive changes, as well as for individual differences.
Taken together, abstract linguistic theories alone are insufficient for establishing direct links between language and neural mechanisms. Bridging this gap requires translating linguistic constructs into computable models and testing them quantitatively against neural data. Computational neuroscience has emerged in this context as a key platform for connecting linguistic theory with brain mechanisms. By integrating neural data, computational modeling, and large-scale language processing models, this field is promoting a shift in language research from descriptive accounts toward mechanistic explanations, and it provides a critical foundation for subsequent discussions of brain-inspired models and advances in language processing.

### v2

@@ Abstract
Elucidating the language–brain relationship requires bridging the methodological gap between linguistics’ abstract theoretical frameworks and neuroscience’s empirical neural data. As an interdisciplinary cornerstone, computational neuroscience formalizes language’s hierarchical and dynamic structures into testable neural representation models through modeling, simulation, and data analysis, enabling computational dialogue between linguistic hypotheses and neural mechanisms. Recent advances in deep learning, particularly large language models (LLMs), have further advanced this inquiry: their high-dimensional representational spaces provide a new scale for probing the neural basis of linguistic processing, while the model–brain alignment framework offers a principled approach to evaluating the biological plausibility of language-related theories.
This review synthesizes interdisciplinary progress from a computational neuroscience perspective. First, it outlines the core connotations of major linguistic frameworks (generative grammar, functional linguistics, and cognitive linguistics), their cross-cultural and evolutionary characteristics, and key challenges for neural alignment, including limited quantitative mechanisms, poor accessibility of abstract constructs to neural measures, and insufficient treatment of dynamics and plasticity. Second, it introduces the methodological foundations of linguistics–neuroscience dialogue, focusing on four technical pillars: neural activity measurement (e.g., fMRI, EEG, MEG, fNIRS, ECoG, SEEG), linguistic numerical representation, the evolution of language models from statistical approaches to LLMs, and neural coding frameworks that link model representations to brain signals, illustrated with a model–brain alignment case study. Third, it summarizes major findings, ranging from early computational insights into predictability and structural processing to recent LLM-driven progress in cross-modal interaction, inter-brain coupling, hierarchical computation, learning strategy sensitivity, and language plasticity. Finally, the review discusses current limitations—including functional alignment without structural homology, constraints on real-time validation, biased research coverage, and narrow evaluation metrics—and proposes future directions, such as improving biological plausibility via spiking neural network–based language models, developing cognitive-level alignment frameworks integrating memory, causality, and metacognition, and extending clinical applications.
@@ 1 Introduction
Recent advances in artificial intelligence, particularly in deep learning and Large Language Models (LLMs), have accelerated this integration further. Researchers can now explore linguistic structures and their correspondence with brain networks within higher-dimensional representational spaces [226, 214, 247]. LLMs provide computational platforms for examining modern linguistic phenomena such as semantic integration, long-distance dependencies, and predictive processing. They also support the emerging framework of model–brain alignment, wherein internal model representations are used to predict and explain neural responses during language processing [56, 5, 117]. This development marks a shift from early conceptual modeling toward quantitative mappings between large-scale linguistic data and high-resolution neural signals.
This paper provides a systematic review of how computational neuroscience serves as a methodological bridge between linguistics and neuroscience. Section 2 surveys modern linguistic frameworks and their development across languages, identifying the core challenges that have historically hindered their alignment with neural data. Section 3 introduces key methodological foundations, including neural data acquisition, word embeddings, modern language models, and neural coding frameworks. Section 4 reviews applications of classical computational neuroscience models to language comprehension, followed by a synthesis of recent progress in LLM-driven neural alignment, covering domains including cross-modal representation, cross-brain coupling, hierarchical computation, and learning mechanisms. Section 5 analyzes current methodological limitations and theoretical challenges, and outlines promising future research directions, such as enhancing biological plausibility, developing cognitive-level alignment frameworks, and expanding clinical and brain–computer interface applications. Finally, Section 6 summarizes the main conclusions and discusses the long-term significance of computational neuroscience for fostering a deeper integration of linguistics and neuroscience.
@@ 2 Theoretical Foundations of Linguistics
Language has long been the core research object defining the intersection of linguistics and neuroscience. Over time, linguistic theories have continued to develop through modeling, cross-cultural research, and empirical investigation. Yet they have also gradually revealed structural limitations that hinder their direct alignment with the neural mechanisms of the brain. Modern linguistic theories offer clear, organized ways to describe the structure and function of language. But when faced with the diversity of languages across cultures, the dynamic nature of everyday language use, and growing evidence from brain science, their ability to fully explain language remains limited [70, 60, 162, 85]. In this section, we review major linguistic perspectives and their evolution across cultural contexts, identifying the key challenges that motivate interdisciplinary integration. This review provides the theoretical background for later discussions on computational neuroscience as a bridging framework, as outlined in Fig. 1.
@@
Figure 1: Modern Linguistic Theories and Challenges for Bridging Linguistics and Neuroscience.
2.1 Frameworks of Modern Linguistic Theories
@@
Modern linguistics comprises several influential theoretical frameworks. Among the most influential are generative grammar, functional linguistics, and cognitive linguistics, each offering a distinct perspective on language structure and use.
@@
Generative grammar, most prominently associated with Noam Chomsky, posits linguistic competence as an innate, biologically grounded component of the human cognitive system [35]. This framework is founded on the hypothesis of a Universal Grammar, arguing that children’s acquisition of complex grammar must be guided by innate principles, given the impoverished and limited nature of their linguistic input. This "poverty of the stimulus" argument is taken as primary evidence for internally constrained, domain-specific grammatical principles. The distinction between deep structure (abstract syntactic relations) and surface structure (observable linguistic forms) illustrates how underlying representations are transformed into observable linguistic expressions [36, 109]. Later developments, most notably the Minimalist Program, sought to reduce linguistic theory to a minimal set of core computational operations, with Merge as the fundamental process, to enhance its cognitive and biological plausibility [134]. Nevertheless, these highly abstract, formal constructs have proven difficult to map onto specific, observable patterns of neural activity, which remains a source of sustained debate in the neurobiology of language.
@@
In contrast, functional linguistics treats language primarily as a system for social communication. Its central tenet is that linguistic structure is shaped and constrained by communicative needs and contextual factors [237]. A major framework within this tradition, Systemic Functional Linguistics, posits that language simultaneously fulfills three core metafunctions: the ideational (for construing experience), the interpersonal (for enacting social relations), and the textual (for organizing discourse). Within this framework, linguistic choices are fundamentally motivated by a speaker’s communicative goals and the dynamics of social interaction [89]. Structural differences between language modalities, such as the prevalence of ellipsis in spoken conversation versus the structural completeness expected in formal writing, are analyzed as adaptive outcomes of the distinct functional demands of each modality [65]. While functional linguistics has been highly influential in discourse analysis, sociolinguistics, and cross-cultural communication studies, its strong emphasis on usage, context, and meaning poses significant challenges for strict formalization. This inherent difficulty with formal modeling, in turn, complicates the establishment of direct, mechanistic links between functional explanations and underlying neural processes.
@@
Cognitive linguistics explicitly rejects the generative conception of language as an autonomous, innate cognitive module. Instead, it argues that language is grounded in, and emerges from, general cognitive processes such as categorization, metaphorical mapping, and mental simulation [130]. Within this framework, Conceptual Metaphor Theory proposes that abstract domains (e.g., time, mind) are understood through systematic mappings from more concrete, embodied experiences, and that grammatical structures often reflect these underlying conceptual representations [122]. Embodied cognition accounts extend this view, proposing that language comprehension is inherently tied to the reactivation of sensorimotor experiences and dynamic mental simulation [9]. While this perspective offers a powerfully unified account of language and cognition, many of its core constructs, such as conceptual metaphors or mental simulations, are difficult to operationalize and quantify experimentally. This lack of form specification consequently limits the framework’s ability to generate precise, testable predictions that can be directly compared with neural data.
@@
Overall, these three major theoretical frameworks account for language in terms of innate mechanisms, communicative function, and general cognition, respectively. However, their explanations remain largely abstract. A substantial gap persists between high-level linguistic descriptions and the concrete neural implementation of language. This gap poses a major challenge for both computational modeling and neurobiological investigation.
@@ 2.2 Cross-Cultural Evolution of Linguistic Theories
The existence of thousands of languages worldwide exhibits remarkable diversity in linguistic structure and expression. This diversity has continuously driven the revision and expansion of linguistic theory, while also underscoring the challenge of constructing a unified explanatory framework, raising questions about whether genuinely universal linguistic features can be abstracted at all [171, 47].
@@
In functional linguistics, cross-linguistic variation is examined in relation to the communicative and cultural contexts in which language is used. Honorific systems, context-dependent expressions, and discourse-level differences are treated not as peripheral but as core objects of grammatical analysis [89]. For example, cultures differ in their emphasis on politeness, indirectness, and social hierarchy, which in turn shape linguistic forms [10]. Although this perspective helps explain cultural variation, its strong reliance on specific cultural contexts makes it difficult to extract universally applicable principles.
Cognitive linguistics examines the interaction between universal embodied experience and culturally specific conceptualizations [130]. Many spatial and emotional metaphors are shared across languages (e.g., time is conceived along a front-back axis, and emotion is metaphorically mapped onto a vertical axis, with happiness represented as "up" and sadness as "down"), suggesting shared cognitive constraints. At the same time, domains such as spatial reference (with systems based on body coordinates like left/right/front/back or on cardinal directions like north/south/east/west), color categorization (with the number and boundaries of basic colour terms varying across languages), and kinship terminology (with varying distinctions between nuclear and extended family) exhibit clear cultural specificity [121]. The coexistence of universality and diversity indicates that language is shaped by both biological constraints and cultural experience.
Cross-linguistic research thus demonstrates that language structure reflects shared cognitive foundations as well as cultural history [39]. However, this twofold origin poses a challenge for determining which linguistic properties stem from universal neural mechanisms and which are products of cultural transmission. Evidence from cross-script studies of developmental dyslexia further illustrates this complexity: reading impairments in alphabetic versus logographic systems engage partially distinct neural circuits, suggesting that the neural correlates of dyslexia reflect both universal constraints and writing-system-specific adaptations [220]. Such nuanced findings underscore the need for closer integration between theoretical linguistics and neurobiologically grounded models of language.
2.3 Challenges in Modern Linguistic Theories
With advances in neuroimaging techniques and computational methods, language research has increasingly moved beyond purely theoretical analysis toward empirical integration with neural data [87]. In this interdisciplinary context, major linguistic theories have revealed several structural challenges. These challenges limit direct links between linguistic constructs and neural mechanisms while motivating the development of new computational and neural modeling approaches [238]. Broadly speaking, these difficulties include a lack of quantitative mechanisms, limited research scope, and poor integrability of findings.
@@ 2.3 Challenges in Modern Linguistic Theories
    Several interpretations have been proposed in the cognitive science literature. Some link the N400 to lexical predictability, suggesting that more predictable words elicit smaller responses. Others interpret it as reflecting semantic integration difficulty [157]. Until recently, these accounts remained largely conceptual, lacking specification of how prediction and integration are implemented computationally within neural circuits. While recent computational models have begun to address this gap, no unifying theory has yet achieved consensus, and the challenge of precise theory-data correspondence remains central to the field.
@@
    Language research often relies on highly controlled experimental designs to ensure reliability and interpretability [240]. Such designs typically focus on decontextualized or localized linguistic phenomena using tightly constrained, uniform stimuli. As a result, findings may not generalize to the richness and variability of natural language use.
    A further challenge lies in the long-standing misalignment of research scales between linguistics and neuroscience [49]. Linguistic studies often examine fine-grained structural distinctions, whereas neuroscience research typically focuses on broader functional organization across brain regions. Even in interdisciplinary work, experiments predominantly rely on simplified or prototypical materials, limiting ecological validity and generalizability.
@@
    While controlled experiments are effective at identifying local effects, they also contribute to fragmented findings. Individual studies typically target specific phenomena, yet language itself resists decomposition into a small set of independent units. Differences in task design, stimulus materials, and measurement techniques further hinder direct integration across studies. For instance, distinct types of linguistic stimuli may elicit divergent activation patterns or electrophysiological responses. However, aggregating such results rarely yields a coherent account of how the brain incrementally constructs the meaning of a complete sentence [99].
@@
    Many central linguistic concepts, such as recursion, thematic roles, and conceptual metaphor, are highly abstract and lack clear operational definitions, making them difficult to translate into measurable neural variables [193]. Moreover, neural activity is distributed across large-scale networks rather than confined to single regions, further complicating attempts to establish simple, one-to-one mappings between linguistic functions and specific neural indicators [62].
@@
    There is no consensus on whether cross-linguistic structural differences arise primarily from innate mechanisms or from cultural learning and experience. These divergent theoretical positions make it difficult for neuroimaging studies to determine whether observed processing differences reflect biological predispositions or experiential adaptation. This uncertainty hinders the development of a unified neurocognitive model of language.
@@
    Major linguistic theories are largely based on static sentence analysis and struggle to account for the dynamic aspects of real-world communication, including real-time prediction, contextual updating, and social interaction processes under naturalistic conditions [189, 188]. In addition, human language abilities show substantial plasticity during child acquisition, second language learning, and recovery after brain injury. These theoretical frameworks lack a unified account of these developmental and adaptive changes, as well as of individual differences in language processing.
Taken together, abstract linguistic theories alone are insufficient for establishing direct links between language and neural mechanisms. Bridging this gap requires translating linguistic constructs into computable models and testing them quantitatively against neural data. Computational neuroscience has emerged in this context as a key platform for connecting linguistic theory with brain mechanisms. By integrating neural data, computational modeling, and large-scale models of language processing, this field is promoting a shift in language research from descriptive accounts toward mechanistic explanations and laying a critical foundation for subsequent work on brain-inspired models and advances in language processing.